OpenStudy (sloppycanada):

How do I find the exact value of 2cos(x) when they give me something for sine?

Nnesha (nnesha):

what do you mean *something for sine* ?

Nnesha (nnesha):

like fraction ? (opposite /hypotenuse ) values ?

OpenStudy (sloppycanada):

like \[\frac{ \sqrt5 }{ 5 }\]

Nnesha (nnesha):

hmm if \[\large \rm \sin(x) =\frac{\sqrt{5}}{5}\] according to cos definition \[\rm sin \rm \theta = \frac{ opposite }{ hypotenuse }~~~~ \cos \theta = \frac{ adjacent }{ hypotenuse } ~~\tan \theta = \frac{ opposite }{ adjacent }\] draw a right triangle find the cos x and then multiply by 2 that's it i guess

OpenStudy (sloppycanada):

It asks me and I quote "Find the exact value of cos2x given that sine = sqrt5/5

Nnesha (nnesha):

cos (2x) or 2 cos(x) or cos^2(x) which one is correct?

OpenStudy (sloppycanada):

cos2x sorry.

OpenStudy (anonymous):

\[2 \cos \left(\sin ^{-1}\left(\frac{1}{\sqrt{5}}\right)\right)=\frac{4}{\sqrt{5}} \]

OpenStudy (sloppycanada):

\[\frac{ \sqrt5 }{ 5 }\] How did you get that?

Nnesha (nnesha):

how did i get sqrt{5}/5 ??

Nnesha (nnesha):

\[\rm \cos (2x) ~~or~~ \cos^2(x)\] which one is correct ?? 2 is square or is it cos(2x) then you can apply double angle formula..

OpenStudy (anonymous):

The sine of x is given as sqrt{5}/5 .

OpenStudy (sloppycanada):

This is what it says - http://prntscr.com/9q9hjt

Nnesha (nnesha):

alright then sqrt{5}/5 is given. as mentioned above.

OpenStudy (mathmale):

I'd use the Pythagorean Theorem. Just suppose you were given sin x = a/h. h would be the hypotenuse, and a the opporite side. The adjacent side (which you need for writing the cosine of the angle x) would be \[\pm \sqrt{h^2-opp^2}.\]

OpenStudy (mathmale):

then the cosine of x would be \[\pm \frac{ adj.side }{ hypotenuse}\]

OpenStudy (mathmale):

Or, use the identity\[\sin^2\theta + \cos^2\theta =1 \] to find cos theta.

Nnesha (nnesha):

its ga ood idea to draw a right triangle. use the given information at which quadrant should it located (this information needed to find is cos negative or positive )  and ofc definitely need Pythagorean theorem a^2+b^2=c^2 where c = hypotenuse(longest side of right triangle )